OpenStudy (anonymous):

Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = _______

OpenStudy (anonymous):

@ranga?

OpenStudy (anonymous):

They don't give f(x) ?

OpenStudy (anonymous):

it's the second line

OpenStudy (ranga):

Looking at the graph you can see g(x) is f(x) shifted up by 4 units. So what should k be in g(x) = f(x) + k?

OpenStudy (anonymous):

i don't really know. Sorry i'm bad at math

OpenStudy (anonymous):

k is the vertical translation of g(x) take a look at the y-int of g(x) and y-int of f(x) calculate the distance between this intercepts and that will be the value of k.

OpenStudy (anonymous):

you're in high school, aren't you ?

OpenStudy (anonymous):

can you help me on another question? And I'm home schooled A State Park charges $5.00 per car plus $1.00 per person as an admission fee. The total charged for a car with x people is f(x) = x + 5. How will the graph of this function change if the per car charge is changed to $8.00 per car? The line will shift vertically down by $3.00. The line will shift vertically up by $3.00. The line will shift vertically up by $8.00. The line will shift vertically down by $8.00.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

same thing as you did from the graph.. since the value of per car is 5 then you will RAISE to 8, what will be the difference between these prices?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

yeah (^_^)

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

i'll try though

OpenStudy (anonymous):

The graph shows the temperature of ice cream in an ice chest during an overnight trip. What is the domain of this function? Graph showing a line segment that goes from ordered pairs 0, negative 10 to 30, 30. The domain is negative 10 through 30. The domain is all integers from negative 10 through 30. The domain is all real numbers 0 through 30. The domain is all integers from 0 through 30.

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

???

OpenStudy (anonymous):

negative 10 to 30, because these were the graphs starts and stops

OpenStudy (anonymous):

okay @Data_LG2 can you help me with a few more? A student is running a 5-kilometer race. He runs 1 kilometer every 3 minutes. Select the function that describes his distance from the finish line. f(x) = - 1 over 3 x + 5 f(x) = - 1 over 5 x + 3 f(x) = 1 over 5 x + 5 f(x) = 1 over 3 x + 3

OpenStudy (anonymous):

lol, someone just asked this question a while ago XD is this a quiz or something? check this: http://openstudy.com/users/data_lg2#/updates/5271781de4b077aaee83cfb1

OpenStudy (anonymous):

sorry i don't see it

OpenStudy (anonymous):

@Data_LG2 can you help me with one last problem? A school puts on a play. The play costs $1,200 in expenses. The students charge $4.00 for tickets. There will be one performance of the play in an auditorium that seats 500 people. What is the domain of the function that shows the profit as a function of the number of tickets sold? The domain is all real numbers from negative 1,200 to positive 2,000 The domain is all integers from negative 1,200 to positive 2,000. The domain is the integers from 0 to 500. The domain is all real numbers from 0 to 500.

OpenStudy (anonymous):

@warood? please help me

OpenStudy (anonymous):

T_T its okay u'll figure i out buddy.

OpenStudy (anonymous):

but im pretty sure you did